OpenStudy (bwaybjorkegren):

Demonstrate how you can use the slope formula to write the point-slope form of an equation of a line.

OpenStudy (shawn):

suppose y = mx + b, and suppose (xₒ,yₒ) is on the line. Then, yₒ = mxₒ + b subtract to get y - yₒ = mx - mxₒ y - yₒ = m(x - xₒ), which is point-slope form

OpenStudy (skullpatrol):

But you @shawn haven't used the slope formula as the question asks?

OpenStudy (shawn):

??? isn't slope formula y = mx + b?

OpenStudy (skullpatrol):

nope, that is the slope-intercept form of an equation of a line.

OpenStudy (shawn):

so what is the slope formula? ax + by + c = 0?

OpenStudy (skullpatrol):

$$\huge m = \dfrac{y_2 - y_1}{x_2 - x_1}$$

OpenStudy (skullpatrol):

That^ is the formula for the slope of a line.

OpenStudy (shawn):

ohhh..... To the asker: in which case just replace y2 with y and x2 with y, and solve for y - y1

OpenStudy (skullpatrol):

Provided $$\huge x_1 \neq x_2$$

OpenStudy (skullpatrol):

Point slope form: $$\huge y-y_0=m(x-x_0)$$

OpenStudy (skullpatrol):

So multiply both side of the slope formula by $$x_2-x_1$$

OpenStudy (skullpatrol):

note you now have to change the subscripts too ;-)